### v2

# CP-LLM: Context and Pixel Aware Large Language Model for Video Quality Assessment

Abstract
Video quality assessment (VQA) is a challenging research topic with broad applications. Effective VQA necessitates sensitivity to pixel-level distortions and a comprehensive understanding of video context to accurately determine the perceptual impact of distortions. Traditional hand-crafted and learning-based VQA models mainly focus on pixel-level distortions and lack contextual understanding, while recent LLM-based models struggle with sensitivity to small distortions or handle quality scoring and description as separate tasks. To address these shortcomings, we introduce CP-LLM: a Context and Pixel aware Large Language Model. CP-LLM is a novel multimodal LLM architecture featuring dual vision encoders designed to independently analyze perceptual quality at both high-level (video context) and low-level (pixel distortion) granularity, along with a language decoder subsequently reasons about the interplay between these aspects. This design enables CP-LLM to simultaneously produce robust quality scores and interpretable quality descriptions, with enhanced sensitivity to pixel distortions (e.g.compression artifacts). The model is trained via a multi-task pipeline optimizing for score prediction, description generation, and pairwise comparisons. Experiment results demonstrate that CP-LLM achieves state-of-the-art cross-dataset performance on established VQA benchmarks and superior robustness to pixel distortions, confirming its efficacy for comprehensive and practical video quality assessment in real-world scenarios.
1 Introduction
Video quality assessment (VQA) is a fundamental research area with broad applications in video compression, transcoding, transmission, playback, content search, and recommendation systems. While subjective testing by human raters remains the most reliable method for quality evaluation, its inherent expense, time-consuming nature, and scalability limitations preclude its application at large scale. Consequently, the development of accurate and robust objective VQA models has been a central focus of research over the past decades. Objective VQA models can be broadly categorized based on the availability of a reference video: no-reference models assess a single target video, whereas full-reference models compare the target video against a pristine reference wang2004image ; li2018vmaf ; zhang2018unreasonable . However, full-reference models operate under the assumption that the reference video is of superior quality, an assumption that falters when dealing with non-pristine or enhanced original content, thereby limiting their reliability. Conversely, certain no-reference models wang2021rich have demonstrated considerable potential in applications traditionally dominated by full-reference approaches, such as video compression assessment, while offering greater flexibility for evaluating enhanced videos. Given these considerations, this paper primarily focuses on the advancement of no-reference VQA.
Figure 1: Videos with intentional effects but incorrectly identified as quality defects by traditional VQA models (CoINVQ wang2021rich , FastVQA wu2022fast , and DOVER wu2022disentangling ), while MLLM’s assessment aligns better with human perception. The first two videos contain intentional low-light and background blurring. The third video presents an astronomy video containing complex visual textures. The last video shows a pixel-art style game characterized by inherent blockiness. All quality scores are calibrated to [1,5][1,5][ 1 , 5 ], where 111, 222, 333, 444, 555 correspond to bad, poor, fair, good, and excellent respectively.
Figure 2: An original video (left) and its compressed version (right). The quality degradation is easily detected by human and VQA models (e.g.CoINVQ), but not correctly caught by MLLMs.
Early no-reference VQA models predominantly concentrated on quantifying low-level signal distortions, such as blur, noise, and blocking artifacts saad2014blind ; mittal2015completely ; korhonen2019two ; tu2021ugc ; korhonen2020blind . However, with the emergence of user-generated content (UGC) videos, several subjective studies wang2019youtube ; wang2024youtube have demonstrated that identical low-level artifacts can elicit significantly different perceptual quality ratings. Consequently, effective UGC-VQA necessitates a comprehensive analysis encompassing both low-level pixel distortions and high-level semantic information. Recognizing this imperative, an increasing number of recent methodologies have aimed to explicitly integrate high-level semantic insights into their model architectures. For instance, Wang et al. wang2021rich developed a learning-based model employing three specialized subnetworks to independently address content, distortion, and compression aspects. These subnetworks subsequently feed their extracted representations into an aggregation network for final score regression. Adopting a similar strategy, Tu et al. tu2021rapique combined traditional hand-crafted low-level signals with features derived from a convolutional neural network (CNN) to capture high-level semantics, thereby enhancing quality prediction.
Although approaches mentioned above underscored the value of leveraging semantic signals, they possess limitations, notably in their content understanding capabilities. As illustrated in Figure 1, certain visual artifacts (e.g., dark scenes, intentional blur, pixelation) that are integral to the artistic style or narrative of the content are often misinterpreted as quality defects by state-of-the-art (SOTA) VQA models such as CoINVQ wang2021rich , FastVQA wu2022fast , and DOVER wu2022disentangling . In contrast, advanced multimodal large language models (MLLMs) like Gemini and GPT have demonstrated the capacity to correctly interpret these stylistic choices as intended, thereby rating such examples as high quality and highlighting a crucial capability absent in many current VQA models.
The emergence of MLLMs offers a promising new paradigm for VQA, as explored in recent studies wu2024comprehensive ; wu2024qbench ; zhang2024qbenchvideo . However, current MLLM applications in VQA encounter several critical challenges. Many implementations rely on pre-existing MLLMs whose vision encoders are primarily optimized for high-level semantic understanding rather than detailed perceptual analysis. While this semantic orientation is advantageous for interpreting diverse content (as exemplified in Figure 1), it can impede nuanced quality assessment, particularly in discerning pixel-level distortions (see Figure 2). Furthermore, the vision encoders in MLLMs typically operate on fixed, low-resolution inputs, necessitating input image resizing that frequently disregards the original aspect ratio and can consequently introduce additional localized distortions. Although some recent work you2024depicting utilized a full-reference paradigm to compare images from the same source with varying artifacts, the assessment of distortion levels still remain coarse, failing to guarantee sensitivity to pixel distortions that impact perceived video quality. Additionally, a majority of existing works utilize MLLMs either exclusively for generating quality-related descriptions wu2024qinstruct ; wu2024towards ; you2024depicting ; you2024descriptive or solely for predicting quality scores wu2024qalign ; you2025deqascore ; wen2024ensemble ; ge2024lmm .
To address these prevailing limitations, we introduce CP-LLM, a novel MLLM-based VQA framework explicitly designed to detect pixel-level artifacts while simultaneously providing robust score predictions and pertinent content- and quality-related descriptions. The work delivers the following key contributions:
  • An MLLM-based VQA model capable of simultaneously generating quality score predictions and interpretable perceptual descriptions.
  • A novel framework ensuring the model excels concurrently at single-video quality prediction, pairwise quality comparisons, and the generation of coherent quality descriptions.
  • A comprehensive experiment design that evaluates both accuracy and robustness of VQA model, with novel evaluation criteria (e.g.flip rate) that have not been sufficiently addressed in existing VQA evaluations.
2 Related Work
Knowledge-driven and Learning-based Methods. Objective methods for assessing UGC video quality have evolved significantly in past decades. Early knowledge-driven approaches, such as V-BLIINDS saad2014blind , VIIDEO mittal2015completely , TLVQM korhonen2019two , and VIDEVAL tu2021ugc , primarily relied on handcrafted spatial and temporal features. Subsequently, hybrid models like CNN-TLVQM korhonen2020blind and RAPIQUE tu2021rapique combined these handcrafted features with representations learned by pretrained CNNs. More recently, purely learning-based models become more and more popular. One common strategy is to employ pretrained networks as separate feature extractors and aggregate features by a subsequent regression module, where relevant works includes VSFA li2019quality , Li22 li2022blindly , PVQ ying2021patch , and CoINVQ wang2021rich . Another direction focuses on end-to-end fine-tuning of the entire network, as demonstrated in SimpleVQA sun2022deep and MinimalisticVQA sun2024analysis . Further advancements include specialized end-to-end techniques, such as the fragment-based processing in FastVQA wu2022fast , the incorporation of aesthetic features in DOVER wu2022disentangling , and the use of modular designs targeting specific quality factors like content, resolution, and frame rate in MBVQA wen2024modular .
MLLM-based Methods. Different from traditional VQA models, MLLM-based methods leverage the extensive knowledge embedded within MLLMs to provide detailed descriptions related to video content and enhance the interpretability of quality assessment. One branch of this research focuses primarily on improving the descriptive capabilities of MLLMs for VQA tasks. For instance, Q-Instruct wu2024qinstruct , Co-Instruct wu2024towards , DepictQA you2024depicting , and DepictQA-Wild you2024descriptive employ specialized datasets and tailored fine-tuning strategies to enhance the low-level perceptual abilities required for nuanced quality descriptions. Conversely, other approaches prioritize direct quality score prediction. Models like Q-Align wu2024qalign and DeQA-Score you2025deqascore harness the perceptual understanding of MLLMs specifically to generate accurate scores. Similarly, LLM-VQA ge2024lmm incorporated SlowFast feichtenhofer2019slowfast motion features to improve the capture of temporal distortions, aiming for more accurate score prediction.
3 CP-LLM: Context and Pixel aware Large Language Model
3.1 Model Components
To achieve good context understanding capability and good sensitivity to pixel distortions, the proposed CP-LLM model incorporates an MLLM architecture with dual vision encoders: one encoder is dedicated to extracting high-level semantic information, while the other focuses on capturing low-level, fine-grained texture details. Features derived from both vision encoders are processed by a quality predictor to get an overall quality score. Meanwhile, these features are concatenated with prompt embeddings and fed into a language decoder to generate comprehensive quality descriptions. The overall structure of CP-LLM is depicted in Figure 3, and the details of four major components are discussed in the following subsections.
Figure 3: The overview of CP-LLM framework. The model processes video frames through two distinct visual pathways: a high-level encoder extracts semantic features, and a low-level encoder captures fine-grained perceptual details from image patches. Embeddings from both visual streams are interleaved with text prompt embeddings and fed into a language decoder to generates textual output. Concurrently, both the concatenated visual embeddings are processed by an MLP regression head to predict a numerical quality score.
3.1.1 High-level Encoder
The high-level vision encoder extracts semantic information and supports downsampling input frames to a lower resolution, ensuring a comprehensive understanding of the overall video context.
We denote a video sequence as $\bm{x}=\{\bm{x}_{i}\}_{i=0}^{N-1}$, where $\bm{x}_{i}\in\mathbb{R}^{H\times W\times 3}$ represents the $i$-th frame, $H$ and $W$ are the original frame height and width, and $N$ is the total number of frames in the sequence. We first uniformly sample a sparse set of $M$ key frames, denoted as $\bm{y}=\{\bm{y}_{i}\}_{i=0}^{M-1}$, from the original sequence $\bm{x}$. Each selected key frame $\bm{y}_{i}\in\mathbb{R}^{H\times W\times 3}$ undergoes preprocessing. A vision encoder $\bm{f}_{h}(\cdot)$ processes the input frame $\bm{y}_{i}$ after bilinear resizing to $H_{h}\times W_{h}$. This encoder outputs a sequence of contextualized visual embeddings $\bm{u}_{i}=\bm{f}_{h}(\bm{y}_{i})$. $\bm{u}_{i}$ is then are then projected through a linear layer to align with a target dimension $D$ (specifically, the embedding dimension of language decoder), resulting $\bm{u}_{i}\in\mathbb{R}^{T_{h}\times D}$.
3.1.2 Low-level Encoder
Targeting fine-grained pixel distortions, the low-level vision encoder operates on input frames at a significantly higher resolution and accurately preserves their aspect ratio to prevent introducing unexpected distortions.
The frame $\bm{y}_{i}$ is resized using bilinear interpolation to a fixed resolution $H_{l}\times W_{l}$. Subsequently, we extract $K$ non-overlapping patches, $\mathcal{P}_{i}=\{\bm{p}_{i,j}\}_{j=0}^{K-1}$, uniformly from the resized frame, where each patch $\bm{p}_{i,j}\in\mathbb{R}^{H_{p}\times W_{p}\times 3}$ has a size $H_{p}\times W_{p}$. These $K$ patches for each key frame $\bm{y}_{i}$ are then fed into a vision Transformer (ViT) model dosovitskiy2020image , denoted as $\bm{f}_{l}(\cdot)$. $\bm{f}_{l}$ processes the patch sequence $\mathcal{P}_{i}$ and outputs a corresponding sequence of feature embeddings $\bm{h}_{i}=\{\bm{h}_{i,j}\}_{j=0}^{K-1}$, where $\bm{h}_{i,j}\in\mathbb{R}^{T_{q}\times D_{q}}$, $T_{q}$ is the number of tokens and $D_{q}$ is the feature dimension of $\bm{f}_{l}$. The features $\bm{h}_{i,j}$ are then projected through a linear layer to align with a target dimension $D$ as well. To obtain a fixed-size representation for the frame while reducing the sequence length, we apply average pooling across the patch dimension and followed by a layer normalization, resulting in the final low-level embeddings $\bm{v}_{i}\in\mathbb{R}^{T_{l}\times D}$.
(a) The input pipeline of CP-LLM requires two inputs (case \@slowromancapi@), and a single input frame (case \@slowromancapii@) will be duplicated to serve as padding and then masked out from the decoder.
(b) Visualization of the attention mask for CP-LLM, configured with a prefix language model attention strategy.
Figure 4: The input pipeline and corresponding attention mask of CP-LLM.
3.1.3 Language Decoder
Language decoder is a critical component of the proposed CP-LLM. Unlike most existing VQA models wang2021rich ; wu2022disentangling that rely on a lightweight regressor to fuse high-level and low-level features, a powerful language decoder makes it possible to comprehensively consider all potential interactions between context and pixel distortions, and generate quality descriptions aligned with human reasoning.
To formalize the language decoder process, we denote the input as a pair $(\bm{y}_{i},\bm{t}_{i})$, where $\bm{t}_{i}$ represents the $i$-th input text prompt, typically provided as a sequence of tokens $\bm{t}_{i}=\{\bm{t}_{i,j}\}_{j=0}^{L-1}$, where $L$ is the maximum length of the sequence. The proposed quality-aware vision-language model, which we can represent as a function $\bm{g}(\cdot,\cdot)$, takes the image $\bm{y}_{i}$ and the text $\bm{t}_{i}$ as input, and generates an output text sequence $\bm{z}_{i}=\{z_{i,j}\}_{j=0}^{O-1}$, where $O$ is the maximum length of the generated output.
The visual embeddings derived from the input frame $\bm{y}_{i}$ consist of high level semantic features $\bm{u}_{i}$ and low level perceptual features $\bm{v}_{i}$. Concurrently, the input text prompt tokens $\bm{t}_{i}$ are transformed into text embeddings $\bm{r}_{i}\in\mathbb{R}^{L\times D}$ by the language model’s native embedding layer $\bm{f}_{t}(\cdot)$. The language model decoder $\bm{g}$, then receives an interleaved sequence formed by combining these three sets of features. Specifically, the input sequence for the decoder is constructed by concatenating the two visual embeddings and the embedded text prompt tokens: $[\bm{u}_{i},\bm{v}_{i},\bm{r}_{i}]$. The decoder operates autoregressively to generate the output token sequence $\bm{z}_{i}$. Each output token $\bm{z}_{i,k}$ is predicted based on the comprehensive context provided by the high level embeddings $\bm{u}_{i}$, the low level visual features $\bm{v}_{i}$, and all preceding tokens, which include the input prompt tokens $\bm{t}_{i}$ and previously generated output tokens $\{z_{i,j}\}_{j=0}^{k-1}$. This process is guided by a Prefix-LM attention mechanism, as visualized in Figure 4(b). This generative process models the conditional probability $p(z_{i,k}|\bm{u}_{i},\bm{v}_{i},\bm{t}_{i},z_{i,0},\dots,z_{i,k-1})$, as determined by the decoder $\bm{g}$.
3.1.4 Quality Predictor
To get an overall quality score is a basic requirement for VQA models. However, some recent studies wu2024comprehensive showed that MLLM is not good at directly rating video quality. Also from an efficiency perspective, reasoning through the entire language decoder just for a single quality score sounds a little over-killed. Thus in our CP-LLM framework, we introduce a light weight quality predictor to assess overall quality.
Our video quality predictor, denoted by the function $f(\cdot):\mathbb{R}^{(T_{h}+T_{l})\times D}\mapsto\mathbb{R}$, takes the concatenation of $\bm{u}_{i}$ and $\bm{v}_{i}$ as input and outputs a scalar quality score $q_{i}$. The quality predictor is a regression head, implemented as a two-layer multilayer perceptron (MLP) with a ReLU activation function between the fully connected layers, takes the visual embeddings as input and predicts a per-frame quality score. The overall quality score $q$ for the video sequence $\bm{x}$ is computed as the mean of the quality scores obtained from the $M$ key frames.
3.2 Robust Multi-task Training Pipeline
The proposed CP-LLM model employs a multi-task training pipeline designed to integrate learning from diverse supervisory signals, accepting pairs of video variants as input, as shown in Figure 4(a). This pipeline concurrently optimizes model parameters based on three core objectives central to comprehensive quality assessment:
First, a pairwise quality ranking loss, $\mathcal{L}_{1}$, encourages the model to correctly order the perceived quality between video variants within a pair:
$\mathcal{L}_{1}=\frac{1}{\mathcal{B}}\sum_{i=1}^{\mathcal{B}}\mathop{\text{max}}(0,-(q_{i}^{(1)}-q_{i}^{(2)})(\hat{q}_{i}^{(1)}-\hat{q}_{i}^{(2)})),$ (1)
where $\mathcal{B}$ denotes the batch size, $q_{i}$ represents the ground-truth quality score (either a mean opinion score (MOS) for an original video or a pseudo-MOS for a generated variant), and $\hat{q}_{i}$ is the predicted score. Superscripts (1)(1)( 1 ) and (2)(2)( 2 ) indicate the first and second video in the $i$-th pair, respectively.
Second, a direct quality score regression loss, $\mathcal{L}_{2}$, focuses on accurately predicting the absolute quality score for original videos:
$\mathcal{L}_{2}=\frac{\sum_{i=1}^{\mathcal{B}}\mathbb{I}_{q_{i}^{(1)}\in\mathcal{M}}(q_{i}^{(1)}-\hat{q}_{i}^{(1)})^{2}}{\mathop{\text{max}}(\sum_{i=1}^{\mathcal{B}}\mathbb{I}_{q_{i}^{(1)}\in\mathcal{M}},1)},$ (2)
where $\mathcal{M}$ is the set of ground-truth MOS values for original videos. The indicator function $\mathbb{I}_{q_{i}^{(1)}\in\mathcal{M}}$ ensures $\ell_{2}$ loss is calculated only for pairs where the first video is an original variant.
Third, an autoregressive text generation loss, $\mathcal{L}_{3}$, trains the model to produce relevant natural language descriptions of content and quality attributes:
$\mathcal{L}_{3}=-\frac{1}{\sum_{i=1}^{\mathcal{B}}O_{i}}\sum_{i=1}^{\mathcal{B}}\sum_{k=1}^{O_{i}}\log\hat{p}_{i,k}(z_{i,k}),$ (3)
where $O_{i}$ is the length of the target text sequence $z_{i}$ for the $i$-th sample, and $\hat{p}_{i,k}(z_{i,k})$ is the predicted probability of the $k$-th target token. This standard cross-entropy loss for next-token prediction is applied only to the generated token sequence representing the descriptive output. The total loss function $\mathcal{L}$ is a weighted combination of these individual losses.
4 Experiments
4.1 Experimental Setups
Datasets. The proposed model, CP-LLM, was trained on LSVQ ying2021patch and augmented with compression variants. These variants were generated by encoding the original LSVQ videos using the H.264 codec at 202020 distinct constant rate factor (CRF) levels, ranging from 151515 to 535353. A foundational assumption in this process is that the perceptual quality of a variant encoded at a higher CRF level (i.e., lower bitrate) does not exceed that of a variant encoded at a lower CRF level (i.e., higher bitrate). To train the language decoder component of CP-LLM, videos from LSVQ were annotated using Gemini 2.0, prompted with: “Based on the provided video frames, create a three-sentence summary. First, describe the visual content of the video in detail. Second, identify the style of the video. Third, assess the technical quality.” These generated annotations constitute a crucial part of the training data for the language decoder. For evaluation, the performance of CP-LLM was assessed on five diverse UGC datasets selected for their varied content types and distortion characteristics: YT-Compress wang2019youtube , YouTubeUGC wang2019youtube , LIVE-YT-Gaming yu2022subjective , Shorts-SDR wang2024youtube , and Shorts-HDR2SDR wang2024youtube .
Evaluated VQA models. We compare our proposed model against a representative set of SOTA VQA methods. This includes several prominent deep learning-based models: CoINVQ wang2021rich , FastVQA wu2022fast , and DOVER wu2022disentangling . To benchmark with recent multimodal approaches, we also include the MLLM-based method Q-Align wu2024qalign and one commercial MLLM Gemini 2.0.
Quality Predictor Performance Criteria. The model performance is evaluated by three criteria: the Spearman’s rank correlation coefficient (SRCC) measures the prediction monotonicity, the Pearson linear correlation coefficient (PLCC) assesses the prediction linearity, the flip rate (FR) quantifies sensitivity to pixel-level alterations, measuring the percentage of incorrectly ranked pairs of videos derived from the same source but with different compression levels. Higher values for SRCC and PLCC, and lower values for FR indicate better performance.
Implementation Details. We set the number of key frames $M=5$ per video for LSVQ, Shorts-SDR and Shorts-HDR2SDR, $M=8$ for LIVE-YT-Gaming, $M=19$ for YT-Compress and YouTube-UGC. Key frames were processed via two pipelines: 1) bilinear resizing to $\times 448$ 448 pixels ignoring aspect ratio for the SigLIP chen2023pali encoder, and 2) bilinear resizing preserving aspect ratio to fit within 540×1080540$\times$ 1080540 × 1080 pixels, followed by cropping $K=8$ non-overlapping $\times 224$ 224 pixel patches for the low-level vision encoder VideoPrism zhao2024videoprism . The maximum sequence length was set to $L=512$ tokens. We fine-tuned the Gemma team2024gemma language decoder using low-rank adaptation (LoRA) hu2022lora , applying a rank of $R=4$ to its query, key, and value attention projections. This fine-tuning process utilized the same prompt that guided the initial data annotation phase. During this process, the high-level vision encoder remained fixed, while the entire low-level vision encoder and the quality regression head underwent full fine-tuning. Optimization was performed with the multi-task loss with equal weights using the Adam optimizer kingma2014adam for 500500500 epochs with a batch size of 128128128 and an initial learning rate of $\eta=1\times 10^{-4}$ on a 161616-core TPU v5.
4.2 Experiment Results
[…]
FastVQA wu2022fast 0.759 / 0.768 0.272 / 0.380 0.220
DOVER wu2022disentangling 0.783 / 0.798 0.399 / 0.516 0.170
Q-Align wu2024qalign 0.795 / 0.799 0.507 / 0.581 0.118
CP-LLM (ours) 0.801 / 0.800 0.540 / 0.605 0.085
Table 1: MOS/DMOS correlations and flip rates evaluated on the YT-Compress dataset. The top-222 results are highlighted in bold and underline.
Table 1 compares the VQA model accuracy and robustness on the YT-Compress dataset, evaluating by SRCC/PLCC on mean opinion scores (MOS) and differential mean opinion scores (DMOS), and flip rate (FR) on pairs of compression variants. We can see CP-LLM achieves superior performance on all three dimensions. Specifically, it surpasses Q-Align, another MLLM-based approach. This suggests that CP-LLM not only excels in making accurate quality assessments for single videos but is also sensitive at detecting distortions within the same contents. Moreover, although trained exclusively with H.264 compression, the model demonstrates notable robustness to the VP9 compression present in the YT-Compress dataset.
The cross-dataset MOS correlations are shown in Table 2. These four datasets represent highly challenging and recent VQA benchmarks, with content that diverges substantially from the LSVQ training set. CP-LLM achieves a top-222 ranking across these four datasets, demonstrating satisfactory generalization capabilities.
Table 3 shows the flip rate results observed across 202020 distinct compression variants of the LSVQ-test-1080p dataset, aiming to evaluate the fine-grained distortion sensitivity. CP-LLM showcases superior sensitivity to pixel-level distortions, outperforming all other models. This highlights the efficacy of CP-LLM’s architecture and comprehensive training strategy in capturing pixel-level artifacts. Notably, prominent models such as Gemini 2.0, considered a leading MLLM, and DOVER, a SOTA VQA model, exhibit limitations in reliably distinguishing small pixel-level distortions. This suggests that in the absence of the proposed quality-related pairwise training, even highly capable models may struggle with such fine-grained perceptual distinctions.
In addition to the quantitative results presented above, qualitative examples are provided to further illustrate CP-LLM’s capabilities. Figure 5 demonstrates that when performing inference with a single input, CP-LLM can generate detailed descriptions of content and quality, accompanied by a predicted score. Figure 6 illustrates that during inference with pairwise inputs, CP-LLM is capable of determining which input exhibits superior quality and provides precise quality scores for each input.
[…]
FastVQA wu2022fast 0.730 / 0.747 0.619 / 0.666 0.791 / 0.800 0.530 / 0.652
DOVER wu2022disentangling 0.777 / 0.792 0.668 / 0.726 0.768 / 0.799 0.512 / 0.614
MBVQA wen2024modular 0.788 / 0.804 0.600 / 0.696 0.753 / 0.802 0.559 / 0.686
Q-Align wu2024qalign 0.833 / 0.846 0.611 / 0.687 0.746 / 0.779 0.585 / 0.672
CP-LLM (ours) 0.810 / 0.813 0.620 / 0.702 0.759 / 0.804 0.586 / 0.708
Table 2: Cross-dataset testing of our model against four competing models, all retrained on the official training split of the large-scale LSVQ dataset and tested on other VQA datasets without fine-tuning. The top-222 results on each dataset are highlighted in bold and underline.
[…]
Gemini 2.0 Answer 0.502 0.434 0.385 0.348 0.319 0.181
CoINVQ wang2021rich Score 0.167 0.142 0.123 0.108 0.086 0.006
FastVQA wu2022fast Score 0.256 0.205 0.169 0.150 0.111 0.024
DOVER wu2022disentangling Score 0.268 0.218 0.179 0.156 0.126 0.026
Q-Align wu2024qalign Score 0.162 0.142 0.122 0.105 0.082 0.008
CP-LLM (ours) Score 0.125 0.069 0.048 0.033 0.021 0.002
Table 3: Flip rate test results for our proposed model versus five competing models, evaluated on LSVQ-test-1080p and its compression variants. “Answer / Score” denotes if the model’s comparison output is a textual answer or a numerical score; “Diff” indicates the CRF difference between compression pairs. The top-222 results are highlighted in bold and underline.
4.3 Ablation Studies
We conduct a series of ablation studies to evaluate the impact of key design choices within our proposed model, specifically examining: 1) different input pipeline strategies (Tables LABEL:tab:input_sp and LABEL:tab:input_fr) and 2) various fine-tuning approaches for the high-level vision encoder (Tables LABEL:tab:train_sp and LABEL:tab:train_fr). Each ablated component is assessed based on its performance with MOS from the two of the LSVQ test sets, reported using SRCC/PLCC, and additionally on their compression variants, evaluated using FR.
Analysis of the input pipeline strategy indicates that a mixed approach, combining single-frame and pairwise inputs, yields superior SRCC/PLCC values compared to employing either mode in isolation. While a purely pairwise strategy is most effective for minimizing the FR, the mixed strategy provides a robust overall balance across these performance metrics. Regarding the high-level vision encoder, freezing all layers was found to optimize SRCC/PLCC performance while minimizing the number of trainable parameters and maintaining a competitive FR. Although fine-tuning all layers achieves the lowest absolute FR, freezing all layers still yields satisfactory results for this metric and offers a more favorable trade-off when all performance aspects are considered.
[…]
Single 0.748 / 0.751 0.620 / 0.674
Pairwise 0.667 / 0.684 0.523 / 0.599
Mix 0.851 / 0.850 0.747 / 0.796
[…]
Single 0.217 0.259
Pairwise 0.055 0.063
Mix 0.072 0.082
[…]
Frozen 0.851 / 0.850 0.747 / 0.796
Head 0.800 / 0.802 0.654 / 0.718
All 0.831 / 0.832 0.705 / 0.750
[…]
Frozen 0.072 0.082
Head 0.103 0.122
All 0.055 0.054
Table 4: Ablation experiments.
Figure 5: Single video quality score prediction and textual explanation by CP-LLM.
Figure 6: Pairwise video comparison results by CP-LLM, with zoom-in differences highlighted by the red box (not feed into the model).
5 Conclusion and Discussion
We have presented CP-LLM, a novel MLLM-based framework for VQA that simultaneously predicts quality scores and generates interpretable textual descriptions. Our method addresses key challenges in VQA, particularly the perception of pixel-level distortions and the integration of quantitative and qualitative feedback, leveraging dual vision encoders and a multi-task training strategy. Comprehensive experiments and ablation studies verify the effectiveness of CP-LLM’s design in achieving robust video quality assessment. Additionally, the model’s dual-output architecture provides a comprehensive insight into video quality beyond a single score.
The current work serves as an important step towards MLLM-based VQA, yet limitations for future research persist. For example, the language decoder currently imposes an input token limit, which restricts the number of video frames that can be processed simultaneously from very long sequences. Exploring more efficient video tokenization strategies to extend the effective input context length is a compelling direction for future work. Furthermore, future investigations could focus on MLLMs within active learning frameworks, potentially for generating sophisticated pseudo-label and boosting VQA model performance.